if the whole leg has a moment of inertia of 5.10 kgm^2 about the COM and a mass of 10.5 kg determine the moment of inertia about the hip which is 0.45 metres away from the COM.
Answer #1

I = Iem + Mary = 5.1 + 10.5 x 0.45 = 5.1+2.12625 = 7:22.625 kg mr = 7.23 kg my